OpenStudy (anonymous):

Can someone please explain to me how to solve this: Suppose that y and t vary inversely and that t = 1/5 when w = 4. Write a function that models the inverse variation, and find t when w = 9.

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

there may be some typo here you have y, t, w

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/5 = k/4 k = 0.8 t = 0.8/9 that's not right :(

OpenStudy (misty1212):

ok it is \(w\) varies inversely with \(t\) so \[\huge w=\frac{k}{t}\] to find \(k\) multiply \(\frac{1}{5}\times 4\)

OpenStudy (misty1212):

yeah \(k=0.8\) so it is \[w=\frac{0.8}{t}\]

OpenStudy (anonymous):

ohh

OpenStudy (misty1212):

if \[w=9\] then you have \[9=\frac{0.8}{t}\] so \[t=\frac{0.8}{9}\]

OpenStudy (anonymous):

wait how did you figure out how to set it up. I always get confused where the varibles go.

OpenStudy (misty1212):

\(w, t\) are the variables, \(k\) is some number you are supposed to find if \(w\) varies inversely with \(t\) that means \[w=\frac{k}{t}\] and you have to find the number \(k\)

OpenStudy (anonymous):

I don't have that answer in my answer choices

OpenStudy (misty1212):

can you post a screenshot?

OpenStudy (anonymous):

yes one second

OpenStudy (anonymous):

im sorry it's taking awhile!

OpenStudy (misty1212):

no problem

OpenStudy (anonymous):

a. t = 1/5w ; 4/45 b. t = 1/5w ; 1/5 c. t = 1/20w ; 1/80 d. t = 4/5w ; 4/45

OpenStudy (misty1212):

oh i see they want it the other way \(t=\frac{k}{w}\) no problem

OpenStudy (anonymous):

OKay

OpenStudy (misty1212):

\(k\) is still \(\frac{4}{5}\) so \[t=\frac{\frac{4}{5}}{w}\] or ' \[t=\frac{4}{5w}\]

OpenStudy (anonymous):

t = 4/45

OpenStudy (anonymous):

thank you so much! :) @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]